OpenStudy (anonymous):

Can someone please help me simplify this radical expression. 5√320 and explain how you got it. I really need help!!

OpenStudy (anonymous):

i take it that this is \[\sqrt[5]{320}\]right? the fifth root in other words

OpenStudy (whpalmer4):

2*2*2*2*2 = 32, so \[\sqrt[5]{320}=2\sqrt[5]{10}\]

OpenStudy (anonymous):

does it help to know that \(2^5=32\) so you have \[\sqrt[5]{2^5\times 10}=\sqrt[5]{2^5}\times \sqrt[5]{10}=2\sqrt[5]{10}\]

OpenStudy (whpalmer4):

If by some chance the problem was instead \[5\sqrt{320} = 5\sqrt{2*2*2*2*2*2*5} = 5\sqrt{2^6*5} = 5*2^3\sqrt{5} = 40\sqrt{5}\]